OpenStudy (anonymous):

What is the value of x in the equation below? 12 15 20 24

OpenStudy (anonymous):

This is the equation 6(x-8)=72

OpenStudy (anonymous):

One method....distribute the left side... 6(x-8) = 6(x) - 6(8)

OpenStudy (anonymous):